i don't know how to find the slope, x and y intercept of the following equation:
y-6=(x-3)
i was thinking maybe the slope might be 0 but i'm not positive because i've never seen a problem like this.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

i don't know how to find the slope, x and y intercept of the following equation:
y-6=(x-3)
y-6 = x-3
y = x+3
The slope-intercept form is y = mx+b. m is the slope.
In y = x+3, the coefficient of x is 1.
So the slope, m, is 1.
---------------
Also, in y=mx+b, b is the y-intercept, so that is (0,3). (where it crosses the y-axis, x = 0)
------------------
To find the x-intercept, set y = 0
0 = x+3
x = -3
That's the point (-3,0)

Answer by stanbon(75887)

find the slope, x and y intercept of the following equation:
y-6=(x-3)
-------------------
Write it in the slope-intercept form: y = mx + b where m is the slope.
y = x-3+6
y = x + 3
---------------------
slope = 1
----------------
x-intercept ?:
Let y = 0, then x = -3
x-intercept: (-3,0)
---------------------
y-intercept ?:
Let x = 0, then y = 3
y-intercept: (0,3)
===========================
Cheers,
Stan H.